OpenStudy (anonymous):

get a medal help me 4(6 + 7) = 4 • 6 + x • 7

OpenStudy (anonymous):

4(6 + 7) = 4 • 6 + x • 7

OpenStudy (anonymous):

4(6 + 7) = 4 • 6 + x • 7

OpenStudy (anonymous):

b

OpenStudy (anonymous):

What is the value of x? 4(6 + 7) = 4 • 6 + x • 7 x =

OpenStudy (anonymous):

4(6 + 7) = 4 • 6 + x • 7 use distributive property 4*6 + 4*7 = 4*6+x*7

OpenStudy (anonymous):

What is the value of x?

OpenStudy (anonymous):

Then, 4*6 + 7x

OpenStudy (anonymous):

It's extremely simple, you have to at least try..

OpenStudy (anonymous):

24 + 7x

OpenStudy (anonymous):

31

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4*6 + 4*7 = 4*6+x*7 | | | | 24 + 28 = 24 + 7x

OpenStudy (anonymous):

83?

OpenStudy (anonymous):

You can't add a number with a variable attached to it. You'd have to isolate it.. Add 24 & 28. Then subtract 24 from the right side to the left. So it should look like this now. 28 = 7x

OpenStudy (anonymous):

No no, 24+28 = 52. So, 52 = 24 + 7x

OpenStudy (anonymous):

Then subtract the 24. 52-24 = 24-24 + 7x 28= 7x

OpenStudy (anonymous):

Then, isolate the variable by dividing 7 by both sides. 28/7 = 7x/7 ? = x

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

so x=4

OpenStudy (anonymous):

@K-K-K-Kiaa

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

k thx